OpenStudy (anonymous):

6x - 12 over 10 my answer: 6(x-1) over 5

OpenStudy (anonymous):

6x-12/10 = 6x-6/5 since the division is a higher priority than subtraction. If the order of the operations were changed by use of parentheses, as in (6x-12)/10 then (6x-12)/10 = (3x-6)/5 = 3(x-2)/5

OpenStudy (anonymous):

i'll write down the steps...thank you so so so so much.....